A 65 W electric blanket runs at 24 V .

What is the resistance of the wire in the blanket?

How much current does the wire carry?
Answer #1

here

R = V^2 / P

R = 24 * 24 / 65

R = 8.86 ohm

then

I = P / V

I = 65 / 24

I = 2.7 amp